Experience has shown that a certain lie detector will show a positive reading (indicates a lie) 10% of the time when a person is telling the truth and 95% of the time when a person is lying. Suppose that a random sample of 5 suspects is subjected to a lie detector test regarding a recent one-person crime. Find the probability of observing no positive reading if all suspects plead innocent and are telling the truth.


1
Expert's answer
2021-07-06T10:20:13-0400

"P(\frac{Test \;negative}{Truth}) = 1 -P(\frac{Test \;positive}{Truth}) \\

= 1 -0.10 \\

= 0.90 \\

P(\frac{all\;five\;results\;are\;test\;negative}{All\;suspects \;telling\;truth}) = \\

= p^5 \\

= 0.9^5 \\

= 0.591"